# Crosscap states and duality of Ising field theory in two dimensions

### v1

@@ Abstract
We propose two distinct crosscap states for the two-dimensional (2D) Ising field theory. These two crosscap states, identifying Ising spins or dual spins (domain walls) at antipodal points, are shown to be related via the Kramers-Wannier duality transformation. We derive their Majorana free field representations and extend bosonization techniques to calculate correlation functions of the 2D Ising conformal field theory (CFT) with different crosscap boundaries. We further develop a conformal perturbation theory to calculate the Klein bottle entropy as a universal scaling function [Phys. Rev. Lett. 130, 151602 (2023)] in the 2D Ising field theory. The formalism developed in this work is applicable to many other 2D CFTs perturbed by relevant operators.
@@
Although the two-dimensional (2D) classical Ising model was invented one century ago [1, 2], studies on this model continue to deepen our understanding of critical phenomena at the present time. In the absence of external fields, Kramers-Wannier duality [3] and the exact solution [4, 5, 6, 7, 8] reveal a second-order phase transition at a critical temperature. At this critical point, scale invariance is promoted to conformal invariance [9, 10, 11], identifying the critical theory as the simplest conformal field theory (CFT) in two dimensions, known as the 2D Ising CFT [12]. In the vicinity of the critical point, the correlation length is much larger than the lattice spacing, so low-energy physics can be captured using a continuous field theory description. Considering the two relevant perturbations (thermal and magnetic perturbations, driven by temperature and external field, respectively), the scaling limit of the 2D classical Ising model is described by the 2D Ising field theory [13, 14], which has attracted considerable interest [15, 16, 17, 18, 19, 20].
@@
There have been significant advances in the study of 2D CFTs on non-orientable manifolds [21, 22, 23, 24, 25, 26, 27], such as the Klein bottle and the real projective plane ($\mathbb{RP}^{2}$), but the 2D classical Ising model is relatively less explored in this regard [28, 29, 30, 31]. In general, crosscap boundary states (and the associated crosscap coefficients) are crucial information for understanding the properties of 2D CFTs on non-orientable manifolds. For the 2D Ising CFT, the crosscap coefficients and certain two-point correlation functions on the $\mathbb{RP}^{2}$ manifold have been employed to conduct a nontrivial benchmark in the bootstrap program [32, 33].
@@
In this Letter, we demonstrate that there are at least two distinct crosscap states for the 2D Ising field theory, both at and away from criticality. We begin with a lattice formulation (quantum Ising chain) and propose two physically motivated crosscap states on the lattice. These lattice crosscap states identify Ising spins and dual spins (domain walls) at antipodal points, respectively, and are related to each other through the Kramers-Wannier duality. Remarkably, the overlaps between the lattice crosscap states and the eigenstates of the critical Ising chain are universal (without finite-size corrections), enabling us to derive Majorana free field representations of the crosscap states in the continuum limit. In the context of the 2D Ising CFT, one of these crosscap states is already known [22] but the other has not been discussed in the literature to the best of our knowledge. Away from criticality, we develop a conformal perturbation theory to calculate the overlap of crosscap states with perturbed ground states, which we call crosscap overlap. This formalism is applicable to any 2D CFT perturbed by relevant operators, thus providing a systematic method to calculate the Klein bottle entropy [34, 35, 36, 37, 38, 39] (norm-square of the crosscap overlap) as a universal scaling function of dimensionless coupling strengths [40]. For the 2D Ising field theory, the leading-order expansions of the Klein bottle entropy derived from the conformal perturbation theory are verified in lattice models by numerical simulations. Our findings open up new avenues for exploring many other 2D field theories on non-orientable manifolds.
@@
where $\sigma^{\alpha}_{j}$ ($\alpha=x,z$) are the Pauli operators at site $j$ and $N$ is the total number of sites. We consider even $N$ and adopt periodic boundary condition ($\sigma_{N+1}^{\alpha}=\sigma_{1}^{\alpha}$) throughout this work. The two relevant perturbations in the Ising field theory [Eq. (1)] are realized by adding $H^{\prime}_{\mathrm{latt}}=(1-h_{z})\sum_{j=1}^{N}\sigma_{j}^{z}-h_{x}\sum_{j=1}^{N}\sigma_{j}^{x}$, where the primary field $\varepsilon(x)$ ($\sigma(x)$) is identified as $-\sigma_{j}^{z}$ ($\sigma^{x}_{j}$) with coupling $g_{1}\sim 1-h_{z}$ ($g_{2}\sim h_{x}$).
The Ising chain in Eq. (2) has a global $\mathbb{Z}_{2}$ symmetry, $[H_{\mathrm{latt}},Q]=0$ with $Q=\prod_{j=1}^{N}\sigma_{j}^{z}$. The eigenvalues of $Q$ define $\mathbb{Z}_{2}$ even ($Q=1$) and $\mathbb{Z}_{2}$ odd ($Q=-1$) subspaces, which, following the CFT convention [12], are called Neveu-Schwarz (NS) and Ramond (R) sectors, respectively. If we restrict ourselves to the NS sector, the critical Ising chain (2) is also invariant (self-dual) under the Kramers-Wannier duality transformation. For our purpose, we define the Kramers-Wannier unitary operator as $U_{\mathrm{KW}}=e^{i\frac{\pi}{4}N}\prod_{j=1}^{N-1}(e^{-i\frac{\pi}{4}\sigma_{j}^{z}}e^{-i\frac{\pi}{4}\sigma^{x}_{j}\sigma^{x}_{j+1}})e^{-i\frac{\pi}{4}\sigma_{N}^{z}}$ [41, 42, 43], which acts on lattice operators as $U_{\mathrm{KW}}\sigma_{j}^{z}U^{{\dagger}}_{\mathrm{KW}}=\sigma_{j}^{x}\sigma_{j+1}^{x}$ and $U_{\mathrm{KW}}\sigma_{j}^{x}U^{{\dagger}}_{\mathrm{KW}}=\sigma_{1}^{y}\prod_{l=2}^{j}\sigma_{l}^{z}$ in the NS sector.
@@
[…]
@@
[…]
@@
With the lattice crosscap state proposals in hand, our next task is to identify their field theory counterparts in the continuum limit. To this end, we adopt the strategy of computing the overlaps of $|\mathcal{C}_{\mathrm{latt}}^{\pm}\rangle$ with eigenstates of the critical Ising chain (2). As the eigenstates of the critical Ising chain can be easily identified with primary and descendant states of the 2D Ising CFT, the overlaps would provide the expansion coefficients of the crosscap states in the Ising CFT basis.
@@
A key observation which enables the overlap computation is
@@
$\phantom{=}\;+\frac{1+i}{2}\prod_{j=1}^{N/2}(1-ic_{j}^{\dagger}c_{j+N/2}^{\dagger})|0\rangle_{c}\,,$ (7)
where $|\mathcal{C}^{+}_{\mathrm{latt}}\rangle$ is written as a sum of two fermionic Gaussian states. After Fourier transforming Eq. (7) into momentum space, its overlaps with the eigenstates of Eq. (5) can be calculated analytically. We find that only eigenstates of the form $|\psi_{k_{1}\cdots k_{M}}\rangle=\prod_{\alpha=1}^{M}(id_{-k_{\alpha}}^{\dagger}d_{k_{\alpha}}^{\dagger})|0\rangle_{d}$ ($0<k_{1}<\cdots<k_{M}<\pi$) have nonvanishing overlaps with $|\mathcal{C}^{+}_{\mathrm{latt}}\rangle$ [43]:
@@
$H^{\mathrm{NS}}_{0}=\frac{2\pi}{L}\left[\sum_{n=1}^{\infty}(n-\frac{1}{2})\left(b^{{\dagger}}_{n-\frac{1}{2}}b_{n-\frac{1}{2}}+\bar{b}^{{\dagger}}_{n-\frac{1}{2}}\bar{b}_{n-\frac{1}{2}}\right)-\frac{c}{12}\right]$ (9)
with $b^{{\dagger}}_{n-\frac{1}{2}}$ and $b_{n-\frac{1}{2}}$ being creation and annihilation operators of the left-moving Majorana fermion (right-moving ones are similar). The connection of Eq. (9) with the lattice model is through the identification of low-energy modes in Eq. (5): $(d_{k},d_{k}^{\dagger})\Leftrightarrow(b_{n_{k}-\frac{1}{2}},b^{{\dagger}}_{n_{k}-\frac{1}{2}})$ and $(id_{-k},-id_{-k}^{\dagger})\Leftrightarrow(\bar{b}_{n_{k}-\frac{1}{2}},\bar{b}^{{\dagger}}_{n_{k}-\frac{1}{2}})$. This is valid for $k$ close to ±$\pm\pi$, where the dispersion of the lattice model (5) can be linearized. The Kramers-Wannier duality of Eq. (9) is inherited from the lattice model: $U_{\mathrm{KW}}b_{n-\frac{1}{2}}U^{{\dagger}}_{\mathrm{KW}}=b_{n-\frac{1}{2}}$ and $U_{\mathrm{KW}}\bar{b}_{n-\frac{1}{2}}U^{{\dagger}}_{\mathrm{KW}}=-\bar{b}_{n-\frac{1}{2}}$.
@@
The crosscap Ishibashi states of the 2D Ising CFT are labeled by the primary fields as $|a\rangle\rangle_{\mathcal{C}}$ with $a=\mathbbm{1},\sigma,\varepsilon$ [21, 22]. The crosscap states $|\mathcal{C}_{\pm}\rangle$ obtained in Eq. (10) are just linear combinations of two Ishibashi states
$|\mathcal{C}_{\pm}\rangle=\sqrt{\frac{2+\sqrt{2}}{2}}|\mathbbm{1}\rangle\rangle_{\mathcal{C}}\pm\sqrt{\frac{2-\sqrt{2}}{2}}|\varepsilon\rangle\rangle_{\mathcal{C}}\,,$ (11)
where $|\mathcal{C}_{+}\rangle$ is already known [22] while $|\mathcal{C}_{-}\rangle$ is a new result. Under the Kramers-Wannier duality transformation, the crosscap Ishibashi states transform as $|\mathbbm{1}\rangle\rangle_{\mathcal{C}}\leftrightarrow|\mathbbm{1}\rangle\rangle_{\mathcal{C}}$ and $|\varepsilon\rangle\rangle_{\mathcal{C}}\leftrightarrow-|\varepsilon\rangle\rangle_{\mathcal{C}}$. Thus, like their lattice counterparts [Eqs. (3) and (4)], two crosscap states in the continuum [Eq. (10) or (11)] are also related through the Kramers-Wannier duality.
Crosscap correlators — The two crosscap states $|\mathcal{C}_{\pm}\rangle$ are indistinguishable from the partition function (with crosscap boundaries): $\langle\mathcal{C}_{+}|e^{-\beta H_{0}}|\mathcal{C}_{+}\rangle=\langle\mathcal{C}_{-}|e^{-\beta H_{0}}|\mathcal{C}_{-}\rangle$. To distinguish them, we calculate the crosscap correlators, which are defined as the conformal correlation functions on the semi-infinite cylinder with a crosscap state at the boundary, in the complex coordinate $z=\tau+ix$ ($\tau\in[0,\infty)$ and $x\in[0,L)$). Geometrically, it is more convenient to perform a conformal transformation $w=e^{\frac{2\pi z}{L}}$ and interpret the crosscap correlators as the correlation functions on the real projective plane ($\mathbb{RP}^{2}$), as depicted in Fig. 1(b).
@@
where $w_{1}$ and $w_{2}$ are the complex coordinates on $\mathbb{RP}^{2}$ and $\eta=\frac{|w_{1}-w_{2}|^{2}}{(1+|w_{1}|^{2})(1+|w_{2}|^{2})}$ is the crosscap cross ratio. The correlator with the boundary state $|\mathcal{C}_{+}\rangle$ is indeed consistent with that obtained using the sewing constraints [22] and the conformal partial wave decomposition [32, 33]. Within the bosonization framework, we also obtain multi-point crosscap correlators for the 2D Ising CFT (see Ref. [43]).
Certain crosscap correlators with two different crosscap states $|\mathcal{C}_{\pm}\rangle$ are related to each other. Under the Kramers-Wannier duality transformation $U_{\mathrm{KW}}$, the fields transform as $\varepsilon\leftrightarrow-\varepsilon$ and $\sigma\leftrightarrow\mu$ ($\mu$: disorder field), whereas the vacuum is unchanged, $U_{\mathrm{KW}}|0\rangle_{\mathrm{NS}}=|0\rangle_{\mathrm{NS}}$. Consequently, certain crosscap correlators are related their dual partners, e.g., ${}_{\mathrm{NS}}\langle 0|\sigma(w_{1},\bar{w}_{1})\sigma(w_{2},\bar{w}_{2})|\mathcal{C}_{\pm}\rangle={}_{\mathrm{NS}}\langle 0|\mu(w_{1},\bar{w}_{1})\mu(w_{2},\bar{w}_{2})|\mathcal{C}_{\mp}\rangle$. However, Eq. (12) reveals that the crosscap correlator of the spin field is not equal to that of its dual field, ${}_{\mathrm{NS}}\langle 0|\sigma(w_{1},\bar{w}_{1})\sigma(w_{2},\bar{w}_{2})|\mathcal{C}_{\pm}\rangle\neq{}_{\mathrm{NS}}\langle 0|\mu(w_{1},\bar{w}_{1})\mu(w_{2},\bar{w}_{2})|\mathcal{C}_{\pm}\rangle$, even though they are equal on the plane. This behavior of the crosscap correlators reflects the non-orientable nature of $\mathbb{RP}^{2}$.
Universal scaling functions — Although expanded in the Ising CFT basis, the validity of Ising crosscap states in Eq. (10) extends to the 2D Ising field theory when relevant perturbations are present. Specifically, the crosscap overlap (overlap of crosscap states with perturbed ground states) is a universal scaling function of the dimensionless couplings with respect to the perturbations [40]. Despite the great potential of this universal scaling function (e.g., identifying critical theoreis in numerics [53]), there was no systematic method to compute it analytically. We now take the first step and put forward a conformal perturbation theory below for computing the crosscap overlap expanded as coupling constants, which is applicable to general 2D CFTs with relevant perturbations.
We consider the Hamiltonian $H=H_{0}+H_{1}$, where $H_{0}$ is the Hamiltonian for a unitary CFT defined on a circle of length $L$. $H_{1}=-g\int_{0}^{L}\mathrm{d}x\,\varphi(x)$ is the relevant perturbation, where $\varphi$ is a primary operator with conformal dimension $(h,\bar{h})$ (we choose $h=\bar{h}$ for simplicity and $h<1$ so that the perturbation is relevant) and normalized as $\lim_{x\rightarrow\infty}\lim_{L\rightarrow\infty}x^{4h}\langle\varphi(0)\varphi(x)\rangle=1$ (expectation value taken with respect to the CFT vacuum). Here a single relevant perturbation is considered for notational simplicity, and the extension to multiple perturbations is straightforward. The perturbed ground state, denoted by $|\psi_{0}(s)\rangle$, only depends on the dimensionless coupling $s=gL^{2-2h}$ [54]. Denoting the crosscap state as $|\mathcal{C}\rangle$, we aim at computing the crosscap overlap $\langle\psi_{0}(s)|\mathcal{C}\rangle$, which only depends on the dimensionless coupling $s$ and hence is a universal scaling function [40].

### v2

@@ Abstract
We propose two distinct crosscap states for the two-dimensional (2D) Ising field theory. These two crosscap states, identifying Ising spins or dual spins (domain walls) at antipodal points, are shown to be related via the Kramers-Wannier duality transformation. We derive their Majorana free field representations and extend bosonization techniques to calculate correlation functions of the 2D Ising conformal field theory (CFT) with different crosscap boundaries. Away from criticality, we develop a conformal perturbation theory to calculate the Klein bottle entropy (norm-square of the crosscap overlap) as a universal scaling function [Phys. Rev. Lett. 130, 151602 (2023)]. For the Ising field theory, our analytical results support the conjectured monotonicity of the Klein bottle entropy under relevant perturbations. The formalism provides a general framework for studying perturbed 2D CFTs on non-orientable manifolds.
@@
Although the two-dimensional (2D) classical Ising model was invented one century ago [1, 2], studies on this model continue to deepen our understanding of critical phenomena at the present time. In the absence of external fields, Kramers-Wannier duality [3] and the exact solution [4, 5, 6, 7, 8] reveal a second-order phase transition at a critical temperature. At this critical point, scale invariance is promoted to conformal invariance [9, 10, 11], identifying the critical theory as the simplest conformal field theory (CFT) in two dimensions, known as the 2D Ising CFT [12]. In the vicinity of the critical point, the correlation length is much larger than the lattice spacing, so the long-wavelength physics can be captured using a continuous field theory description. Considering the two relevant perturbations (thermal and magnetic perturbations, driven by temperature and external field, respectively), the scaling limit of the 2D classical Ising model is described by the 2D Ising field theory [13, 14], which has attracted considerable interest [15, 16, 17, 18, 19, 20].
@@
There have been significant advances in the study of 2D CFTs on non-orientable manifolds [21, 22, 23, 24, 25, 26, 27, 28], such as the Klein bottle and the real projective plane ($\mathbb{RP}^{2}$), but the 2D classical Ising model is relatively less explored in this regard [29, 30, 31, 32]. In general, crosscap boundary states (and the associated crosscap coefficients) are crucial information for understanding the properties of 2D CFTs on non-orientable manifolds. For the 2D Ising CFT, the crosscap coefficients and certain two-point correlators on the $\mathbb{RP}^{2}$ manifold have been employed to conduct a nontrivial benchmark in the bootstrap program [33, 34].
@@
In this Letter, we demonstrate that the 2D Ising field theory supports two distinct crosscap states, related by Kramers-Wannier duality, which can be explicitly constructed both at and away from criticality. We begin with a lattice formulation (quantum Ising chain) and propose two physically motivated crosscap states on the lattice. These lattice crosscap states identify Ising spins and dual spins (domain walls) at antipodal points, respectively. Remarkably, the overlaps between the lattice crosscap states and the eigenstates of the critical Ising chain are universal (without finite-size corrections), enabling us to derive Majorana free field representations of the crosscap states in the continuum limit. Within the 2D Ising CFT, one of them is the standard Pradisi-Sagnotti-Stanev (PSS) crosscap state [22], while the other corresponds to an equal-weight superposition of the standard PSS crosscap state and a generalized PSS crosscap state labeled by a simple current [23, 35, 36]. Away from criticality, we develop a conformal perturbation theory to calculate the overlap of crosscap states with perturbed ground states, which we call crosscap overlap. This formalism is applicable to any 2D CFT perturbed by relevant operators, thus providing a systematic method to calculate the Klein bottle entropy [37, 38, 39, 40, 41, 42] (norm-square of the crosscap overlap) as a universal scaling function of dimensionless coupling constants [43]. For the $A$-series unitary minimal models, our conformal perturbation theory analysis provides a perturbative proof of the conjectured monotonicity of the Klein bottle entropy [43]. For the 2D Ising field theory, our leading-order analytical results further support this monotonicity, strengthening the conjecture.
@@
where $\sigma^{\alpha}_{j}$ ($\alpha=x,z$) are the Pauli operators at site $j$ and $N$ is the total number of sites. We consider even $N$ and adopt periodic boundary condition ($\sigma_{N+1}^{\alpha}=\sigma_{1}^{\alpha}$) throughout this work. The two relevant perturbations in the Ising field theory [Eq. (1)] are realized by adding $H^{\prime}_{\mathrm{latt}}=(1-h_{z})\sum_{j=1}^{N}\sigma_{j}^{z}-h_{x}\sum_{j=1}^{N}\sigma_{j}^{x}$, where the primary field $\varepsilon(x)$ ($\sigma(x)$) is identified as $\sigma_{j}^{z}$ ($\sigma^{x}_{j}$) with coupling $g_{1}\sim 1-h_{z}$ ($g_{2}\sim h_{x}$).
The Ising chain in Eq. (2) has a global $\mathbb{Z}_{2}$ symmetry, $H_{\mathrm{latt}},Q]=0$ with $Q=\prod_{j=1}^{N}\sigma_{j}^{z}$. The eigenvalues of $Q$ define $\mathbb{Z}_{2}$ even ($Q=1$) and $\mathbb{Z}_{2}$ odd (Q=1Q=-1) subspaces, which, following the CFT convention [12], are called Neveu-Schwarz (NS) and Ramond (R) sectors, respectively. If we restrict ourselves to the NS sector, the critical Ising chain (2) is also invariant (self-dual) under the Kramers-Wannier duality transformation. For our purpose, we define the Kramers-Wannier unitary operator as $U_{\mathrm{KW}}=e^{i\frac{\pi}{4}N}\prod_{j=1}^{N-1}(e^{-i\frac{\pi}{4}\sigma_{j}^{z}}e^{-i\frac{\pi}{4}\sigma^{x}_{j}\sigma^{x}_{j+1}})e^{-i\frac{\pi}{4}\sigma_{N}^{z}}$ [44, 45, 46], which acts on lattice operators as $U_{\mathrm{KW}}\sigma_{j}^{z}U^{{\dagger}}_{\mathrm{KW}}=\sigma_{j}^{x}\sigma_{j+1}^{x}$ and $U_{\mathrm{KW}}\sigma_{j}^{x}U^{{\dagger}}_{\mathrm{KW}}=\sigma_{1}^{y}\prod_{l=2}^{j}\sigma_{l}^{z}$ in the NS sector.
@@
where $|\Uparrow\rangle\equiv|\uparrow_{1}\uparrow_{2}\cdots\uparrow_{N}\rangle$ is the fully polarized state in the $\sigma^{z}$-basis. When placing $|\mathcal{C}^{+}_{\mathrm{latt}}\rangle$ on the 1D circle, maximally entangled pairs $|\uparrow\uparrow\rangle+|\downarrow\downarrow\rangle$ identify Ising spins between two antipodal sites $j$ and $j+N/2$ [47, 48, 49, 50, 51, 52], as depicted in Fig. 1(a).
@@
where $\mu_{j}=\prod_{l=1}^{j}\sigma_{l}^{z}$ is the Ising disorder operator (dual spin). $|\Rightarrow\rangle\equiv|\rightarrow_{1}\rightarrow_{2}\cdots\rightarrow_{N}\rangle$ and $|\Leftarrow\rangle\equiv|\leftarrow_{1}\leftarrow_{2}\cdots\leftarrow_{N}\rangle$ are fully polarized states in the $\sigma^{x}$-basis, i.e., $|\rightarrow\rangle=\frac{1}{\sqrt{2}}(|\uparrow\rangle+|\downarrow\rangle)$ and $|\leftarrow\rangle=\frac{1}{\sqrt{2}}(|\uparrow\rangle-|\downarrow\rangle)$. As $\sigma^{z}$ flips spins in the $\sigma^{x}$-basis, $\mu_{j}\mu_{j+N/2}$ creates two domain walls at antipodal positions on top of $|\Rightarrow\rangle$ or $|\Leftarrow\rangle$, as illustrated in Fig. 1(a). Thus, the lattice crosscap state $|\mathcal{C}^{-}_{\mathrm{latt}}\rangle$ identifies each pair of dual spins (domain walls) at the antipodal sites.
The two lattice crosscap states satisfy $U_{\mathrm{KW}}|\mathcal{C}^{\pm}_{\mathrm{latt}}\rangle=|\mathcal{C}^{\mp}_{\mathrm{latt}}\rangle$, and therefore, are dual to each other under the Kramers-Wannier transformation. Remarkably, the dual state $|\mathcal{C}^{-}_{\mathrm{latt}}\rangle$ can be expressed as [46]
$|\mathcal{C}^{-}_{\mathrm{latt}}\rangle=\frac{1}{\sqrt{2}}(|\mathcal{C}^{+}_{\mathrm{latt}}\rangle+|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle)\,$, (5)
which is an equal weight superposition of $|\mathcal{C}^{+}_{\mathrm{latt}}\rangle$ and another modified crosscap state, defined by
$|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle=\prod_{j=1}^{N/2}(1-\sigma^{x}_{j}\sigma^{x}_{j+N/2})|\Uparrow\rangle\,.$ (6)
The meaning of these lattice crosscap states will become more transparent in the continuum limit, as we shall demonstrate below.
@@
To identify their field theory counterparts, we adopt the strategy of computing the overlaps of $|\mathcal{C}_{\mathrm{latt}}^{\pm}\rangle$ with eigenstates of the critical Ising chain (2). As the eigenstates of the critical Ising chain can be easily identified with primary and descendant states of the 2D Ising CFT, the overlaps yield the expansion coefficients of the crosscap states in the Ising CFT basis.
@@
A key observation enabling the overlap computation is that $|\mathcal{C}^{+}_{\mathrm{latt}}\rangle$ can be expressed as a sum of two fermionic Gaussian states
@@
$\phantom{=}\;+\frac{1+i}{2}\prod_{j=1}^{N/2}(1-ic_{j}^{\dagger}c_{j+N/2}^{\dagger})|0\rangle_{c}\,.$ (9)
After transforming Eq. (9) into momentum space, its overlaps with the eigenstates of Eq. (7) can be calculated analytically. We find that only eigenstates of the form $|\psi_{k_{1}\cdots k_{M}}\rangle=\prod_{\alpha=1}^{M}(id_{-k_{\alpha}}^{\dagger}d_{k_{\alpha}}^{\dagger})|0\rangle_{d}$ ($0<k_{1}<\cdots<k_{M}<\pi$) have nonvanishing overlaps with $|\mathcal{C}^{+}_{\mathrm{latt}}\rangle$ [46]:
@@
$H^{\mathrm{NS}}_{0}=\frac{2\pi}{L}\left[\sum_{n=1}^{\infty}(n-\frac{1}{2})\left(b^{{\dagger}}_{n-\frac{1}{2}}b_{n-\frac{1}{2}}+\bar{b}^{{\dagger}}_{n-\frac{1}{2}}\bar{b}_{n-\frac{1}{2}}\right)-\frac{1}{24}\right]$ (11)
with $b^{{\dagger}}_{n-\frac{1}{2}}$ ($b_{n-\frac{1}{2}}$) being creation (annihilation) operators of the left-moving Majorana fermion (right-moving ones are similar). The connection of Eq. (11) with the lattice model is through the identification of low-energy modes in Eq. (7): $(d_{k},d_{k}^{\dagger})\Leftrightarrow(b_{n_{k}-\frac{1}{2}},b^{{\dagger}}_{n_{k}-\frac{1}{2}})$ and $(id_{-k},-id_{-k}^{\dagger})\Leftrightarrow(\bar{b}_{n_{k}-\frac{1}{2}},\bar{b}^{{\dagger}}_{n_{k}-\frac{1}{2}})$. This is valid for $k$ close to $\pm\pi$, where the dispersion of the lattice model (7) can be linearized. The Kramers-Wannier duality of Eq. (11) is inherited from the lattice model: $U_{\mathrm{KW}}b_{n-\frac{1}{2}}U^{{\dagger}}_{\mathrm{KW}}=b_{n-\frac{1}{2}}$ and $U_{\mathrm{KW}}\bar{b}_{n-\frac{1}{2}}U^{{\dagger}}_{\mathrm{KW}}=-\bar{b}_{n-\frac{1}{2}}$.
@@
The crosscap Ishibashi states of the 2D Ising CFT are labeled by the primary fields as $|a\rangle\rangle_{\mathcal{C}}$ with $a=\mathbbm{1},\sigma,\varepsilon$ [21]. The crosscap states $|\mathcal{C}_{\pm}\rangle$ obtained in Eq. (12) are just linear combinations of two Ishibashi states [46]
$|\mathcal{C}_{\pm}\rangle=\sqrt{\frac{2+\sqrt{2}}{2}}|\mathbbm{1}\rangle\rangle_{\mathcal{C}}\pm\sqrt{\frac{2-\sqrt{2}}{2}}|\varepsilon\rangle\rangle_{\mathcal{C}}\,.$ (13)
Comparing with the PSS crosscap states for the Ising CFT [23, 35], $|\mathcal{C}_{\mathbbm{1}}\rangle=\sqrt{2}\cos\frac{\pi}{8}|\mathbbm{1}\rangle\rangle_{\mathcal{C}}+\sqrt{2}\sin\frac{\pi}{8}|\varepsilon\rangle\rangle_{\mathcal{C}}$ and $|\mathcal{C}_{\varepsilon}\rangle=\sqrt{2}\sin\frac{\pi}{8}|\mathbbm{1}\rangle\rangle_{\mathcal{C}}-\sqrt{2}\cos\frac{\pi}{8}|\varepsilon\rangle\rangle_{\mathcal{C}}$, it becomes clear that $|\mathcal{C}_{+}\rangle=|\mathcal{C}_{\mathbbm{1}}\rangle$ and $|\mathcal{C}_{-}\rangle=\frac{1}{\sqrt{2}}(|\mathcal{C}_{\mathbbm{1}}\rangle+|\mathcal{C}_{\varepsilon}\rangle)$ [36]. Interestingly, Eqs. (5) and (6) provide a perfect lattice correspondence of the latter relation. It becomes evident once we notice that, in the continuum limit, $|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle$ corresponds to $|\mathcal{C}_{\varepsilon}\rangle$ [46], by the same type of calculation that identifies $|\mathcal{C}^{+}_{\mathrm{latt}}\rangle$ with $|\mathcal{C}_{\mathbbm{1}}\rangle$.
Crosscap correlators — For the Ising CFT, the two crosscap states $|\mathcal{C}_{\pm}\rangle$ are indistinguishable from the partition function (with crosscap boundaries): $\langle\mathcal{C}_{+}|e^{-\beta H_{0}}|\mathcal{C}_{+}\rangle=\langle\mathcal{C}_{-}|e^{-\beta H_{0}}|\mathcal{C}_{-}\rangle$. To distinguish them, we calculate the crosscap correlators, which are defined as the conformal correlation functions on the semi-infinite cylinder with a crosscap state at the boundary, in the complex coordinate $z=\tau+ix$ ($\tau\in[0,\infty)$ and $x\in[0,L)$). Geometrically, it is more convenient to perform a conformal transformation $w=e^{\frac{2\pi z}{L}}$ and interpret the crosscap correlators as the correlation functions on the real projective plane ($\mathbb{RP}^{2}$), as depicted in Fig. 1(b).
@@
where $w_{1}$ and $w_{2}$ are the complex coordinates on $\mathbb{RP}^{2}$ and $\eta=\frac{|w_{1}-w_{2}|^{2}}{(1+|w_{1}|^{2})(1+|w_{2}|^{2})}$ is the crosscap cross ratio. The correlator with the boundary state $|\mathcal{C}_{+}\rangle$ is indeed consistent with that presented in Refs. [22, 33, 34]. Within the bosonization framework, we also obtain multi-point crosscap correlators for the 2D Ising CFT [46].
Universal scaling functions — While expressed in the Ising CFT basis, the validity of Ising crosscap states in Eq. (12) extends to the 2D Ising field theory when relevant perturbations are present. The resulting crosscap overlap, defined as the overlap of crosscap states with perturbed ground states, is a universal scaling function of the dimensionless couplings [43]. Despite its great potential for identifying critical theories in numerics, there was no systematic method to compute it analytically. Here we take the first step by developing a conformal perturbation theory that allows for a perturbative expansion of the crosscap overlap in powers of the coupling constants, applicable to general 2D CFTs with perturbations.
We consider the Hamiltonian $H=H_{0}+H_{1}$, where $H_{0}$ is the Hamiltonian of a unitary CFT defined on a circle of length $L$. The perturbation is given by $H_{1}=-g\int_{0}^{L}\mathrm{d}x\,\varphi(x)$, where $\varphi$ is a primary operator of conformal dimension $(h,\bar{h})$ (we take $h=\bar{h}$ for simplicity and $h<1$ so that the perturbation is relevant), normalized as $\lim_{x\rightarrow\infty}\lim_{L\rightarrow\infty}x^{4h}\langle\varphi(0)\varphi(x)\rangle=1$ with the expectation value taken in the CFT vacuum. For notational simplicity we consider a single relevant perturbation; the extension to multiple perturbations is straightforward. The perturbed ground state, denoted as $|\psi_{0}(s)\rangle$, depends only on dimensionless coupling $s=gL^{2-2h}$ [56]. Denoting the crosscap state as $|\mathcal{C}\rangle$, our goal is to compute the crosscap overlap $\langle\psi_{0}(s)|\mathcal{C}\rangle$, which depends solely on $s$ and is thus a universal scaling function [43].